### v2

# Digraphs in which every $t$ vertices have exactly $\lambda$ common out-neighbors

Abstract
We say that a digraph is a $(t,\lambda)$-liking digraph if every $t$ vertices have exactly $\lambda$ common out-neighbors. In 1975, Plesník [Graphs with a homogeneity, 1975. Glasnik Mathematicki 10:9-23] proved that any $(t,1)$-liking digraph is the complete digraph on $t+1$ vertices for each $t\geq 3$. Choi et al. [A digraph version of the Friendship Theorem, 2025. Discrete mathematics, 348(1), 114238] showed that a $(2,1)$-liking digraph is a fancy wheel digraph or a $k$-diregular digraph for some positive integer $k$. In this paper, we extend these results by completely characterizing the $(t,\lambda)$-liking digraphs with $t\geq\lambda+2$ and giving some equivalent conditions for a $(t,\lambda)$-liking digraph being a complete digraph on $t+\lambda$ vertices.
Keywords. Liking digraph; Generalized Liking digraph; Generalized Friendship graph; Complete digraph; Diregular digraph; $t$-design.
2020 Mathematics Subject Classification. 05C20, 05C75
1 Introduction
In this paper, for graph-theoretical terminology and notations not defined, we follow [1] and [2]. Neither graphs nor digraphs in this paper have loops, multiple edges, or multiple arcs.
This paper was motivated by the Friendship Theorem introduced by Erdös et al. [7] in 1966. Using graph-theoretical terminology, the Friendship Theorem can be described as follows: if any pair of vertices in a graph has exactly one common neighbor, then there exists a vertex adjacent to the others. A graph that satisfies the hypothetical part of the Friendship Theorem is called a “friendship graph”, that is, a friendship graph is a graph such that every pair of vertices has exactly one common neighbor. Many variants of a friendship graph have been studied (see [3, 4, 6, 8, 9, 10, 12, 13, 14, 16]).
A digraph version of the Friendship Theorem was introduced and studied as a variant ([5, 13, 14]). A digraph is called a $(t,\lambda)$-liking digraph if every $t$ vertices have exactly $\lambda$ common out-neighbors for some positive integers $t$ and $\lambda$. In 1974, Müller and Pelant [13] studied the case where tournaments are $(t,\lambda)$-liking digraphs and showed the non-existence of such tournament for $t\geq 3$. In 1975, Plesník [14] characterized the $(t,1)$-liking digraphs for each integer $t\geq 3$ by proving that every $(t,1)$-liking digraph with $t\geq 3$ is the complete digraph on $t+1$ vertices, which is a generalization of the result given by Müller and Pelant. He proposed as an open question to take care of the case $t=2$. In 2025, Choi et al. [5] settled the question by completely characterizing the $(2,1)$-liking digraphs.
In this paper, we extend the results of Plesník and Choi et al. by completely characterizing the $(t,\lambda)$-liking digraphs for any positive integers $t$ and $\lambda$ satisfying $t\geq\lambda+2$ as follows.
Theorem 1.1.
If $t\geq\lambda+2$, then the complete digraph on $t+\lambda$ vertices is the only $(t,\lambda)$-liking digraph.
Indeed, we give some equivalent conditions for a $(t,\lambda)$-liking digraph being the complete digraph on $t+\lambda$ vertices.
Theorem 1.2.
Let $D$ be a $(t,\lambda)$-liking digraph for some positive integers $t,\lambda$ with $t\geq 2$. Then the following are equivalent.
  1. (a)
    $D$ is complete on $t+\lambda$ vertices, that is, $D\cong\overleftrightarrow{K}_{t+\lambda}$.
  2. (b)
    $D$ is a $(t-1,\lambda+1)$-liking digraph.
  3. (c)
    $d^{+}(v)=t+\lambda-1$ for each vertex $v$.
Furthermore, (a) is equivalent to the condition (d) if $(t,\lambda)\neq(2,1)$; (e) if $t\geq 3$, where
  1. (d)
    there is a vertex $v$ satisfying $N^{+}(v)=V(D)\setminus\{v\}$;
  2. (e)
    $D$ is diregular.
We will prove the above two theorems in Section 3.
We note that the equivalence of (a) and (e) in Theorem 1.2 implies that if $t\geq 3$, then the complete digraph on $t+\lambda$ vertices is the only diregular $(t,\lambda)$-liking digraph. This reveals a key structural rigidity specific to the directed setting: diregularity alone forces completeness when $t\geq 3$, a phenomenon that breaks down at $t=2$ (see Figure 1).
There is a variant of a friendship graph, called a “generalized friendship graph”. A generalized friendship graph is a graph if any $t$ vertices have exactly $\lambda$ common neighbors for some positive integers $t$ and $\lambda$. It can be said that a $(t,\lambda)$-liking digraph is a digraph version of a generalized friendship graph. Considering this observation, we refer this graph as a $(t,\lambda)$-friendship graph to make the concept clearer. The structure of $(t,\lambda)$-friendship graph is well-known. In [4] and [16], it was shown that a $(t,\lambda)$-friendship graph is the complete graph on $t+\lambda$ vertices if $t\geq 3$. Theorem 1.1 shows that the digraph version of this theorem is true if $t\geq\lambda+2$. Moreover, the equivalence of (a) and (e) in Theorem 1.2 generalizes the theorem for $(t,\lambda)$-friendship graphs since a $(t,\lambda)$-friendship graph becomes a diregular $(t,\lambda)$-liking digraph by replacing each edge with a directed cycle of length two.
2 Preliminaries
This section describes the existing results needed to prove Theorems 1.1 and 1.2 and the results that can be simply derived from them.
Lemma 2.1 (Lemma 6.2 in [14]).
Let $D$ be a $(t,\lambda)$-liking digraph. Then $|V(D)|\geq t+\lambda$ and $\delta^{+}(D)\geq t+\lambda-1$.
Lemma 2.2 (Lemmas 6.4 and 6.6 in [14]).
Let $D$ be a $(t,\lambda)$-liking digraph. Then the following hold.
  1. (1)
    $\sum_{v\in V(D)}{d^{-}(v)\choose t}=\lambda{|V(D)|\choose t}.$
  2. (2)
    For each $v\in V(D)$,
    (${d^{-}(v)\choose t-1}\leq{d^{+}(v)\choose\lambda}.$
Proposition 2.3 (Lemma 6.3 in [14]).
Let $D$ be a $(t,\lambda)$-liking digraph. Then for any $1\leq i<t$ and $\{v_{1},\ldots,v_{i}\}\subseteq V(D)$, the subdigraph of $D$ induced by the common out-neighbors of $v_{1},\ldots,v_{i}$ is a $(t-i,\lambda)$-liking digraph.
We make a meaningful observation from Lemma 2.1 and Proposition 2.3.
Proposition 2.4.
Let $D$ be a $(t,\lambda)$-liking digraph. For each $1\leq i<t$, every $t-i$ vertices have at least $\lambda+i$ common out-neighbors in $D$.
Proof.
Fix $i\in\{1,\ldots,t-1\}$. Take $t-i$ vertices. Then the subdigraph of $D$ induced by the common out-neighbors of the $t-i$ vertices is a $(i,\lambda)$-liking digraph by Proposition 2.3, and so it has at least $\lambda+i$ vertices by Lemma 2.1. Thus every $t-i$ vertices have at least $\lambda+i$ common out-neighbors in $D$. ∎
Given a digraph $D$ and vertices $u$ and $v$ of $D$, we use the expression $u\to v$ when $(u,v)\in A(D)$. When representing negation, add a slash (///) to the symbol.
A $k$-diregular digraph is a digraph in which each vertex has outdegree $k$ and indegree $k$ for a positive integer $k$. A digraph is said to be diregular if it is a $k$-diregular digraph for some positive integer $k$.
Let $v$, $k$, $\lambda$, and $t$ be positive integers such that $v>k\geq t$. A $t$-$(v,k,\lambda)$ design $\mathfrak{D}=(X,\mathcal{B})$ consists of a set $X$ of elements, called varieties, and a collection $\mathcal{B}$ of subsets of $X$, called blocks, such that the following conditions are satisfied:
  • $|X|=v$;
  • Each block consists of exactly the same number $k$ of varieties;
  • Every set of $t$ distinct varieties appears simultaneously in exactly the same number $\lambda$ of blocks.
The general term $t$-design is used to indicate any $t$-$(v,k,\lambda)$ design.
It is a well-known fact that the number of blocks in a $t$-$(v,k,\lambda)$ design is $\lambda{v\choose t}/{k\choose t}$ and each variety appears exactly the same number $\lambda{v-1\choose t-1}/{k-1\choose t-1}$ of blocks. A $t$-$(v,k,\lambda)$ design is said to be symmetric if the number of its blocks is equal to the number of its varieties. That is,
$v=\frac{\lambda{v\choose t}}{{k\choose t}}.$
Then each variety of a symmetric $t$-$(v,k,\lambda)$ appears in exactly $k$ blocks. [15]
Lemma 2.5.
If a $k$-diregular $(t,\lambda)$-liking digraph of order $n$ exists, then there is a symmetric $t$-$(n,k,\lambda)$ design.
Proof.
Suppose that there is a $k$-diregular $(t,\lambda)$-liking digraph $D$ of order $n$. To form a symmetric $t$-$(n,k,\lambda)$ design, we take the vertices of $D$ as varieties and the in-neighborhoods of the vertices as blocks. Since $D$ is $k$-diregular, each block has size $k$. Moreover, since $D$ is a $(t,\lambda)$-liking digraph, each $t$-subset of varieties appears in exactly $\lambda$ blocks. Thus we have obtained a symmetric $t$-$(n,k,\lambda)$ design. ∎
Proposition 2.6 (Lemma 2.6 in [11]).
If a $t$-$(v,k,\lambda)$ design with $t\geq 3$ is symmetric, then $k\geq v-1$.
A digraph $D$ is complete if, for every pair $x$, $y$ of distinct vertices of $D$, both $x\to y$ and $y\to x$. The complete digraph on $n$ vertices is denoted by $\overleftrightarrow{K}_{n}$.
By Lemma 2.5 and Proposition 2.6, the following holds.
Proposition 2.7.
Let $D$ be a diregular $(t,\lambda)$-liking digraph with $t\geq 3$. Then $D$ is the complete digraph on $t+\lambda$ vertices.
Proof.
Suppose that $D$ is $k$-diregular for some positive integer $k$. Then, since $D$ has neither loops nor multiple arcs, $k<|V(D)|$. By Lemma 2.5 and Proposition 2.6, $k\geq|V(D)|-1$. Thus $k=|V(D)|-1$. That is, $N^{+}(v)=V(D)-\{v\}$ for each vertex $v$. Therefore $D$ is complete. Since $D$ is a $(t,\lambda)$-liking digraph, $|V(D)|=t+\lambda$. ∎
3 Proofs of Theorems 1.1 and 1.2
The case where $t=1$ for both Theorems 1.1 and 1.2 is excluded, so we only consider $(t,\lambda)$-liking digraphs with $t\geq 2$ and $\lambda\geq 1$.
Lemma 3.1.
Let $D$ be a $(t,\lambda)$-liking digraph. If $d^{+}(v)\leq d^{-}(v)$ for a vertex $v$, then
$0\leq(d^{+}(v)-t-\lambda+1)(\lambda-t+1).$
Furthermore, if the conditional inequality is strict, then the resulting inequality is also strict.
Proof.
Suppose that a vertex $v$ satisfies
$d^{+}(v)\leq d^{-}(v).$
Then
(${d^{+}(v)\choose t-1}\leq{d^{-}(v)\choose t-1}.$
Thus, by Lemma 2.2(2),
(${d^{+}(v)\choose t-1}\leq{d^{+}(v)\choose\lambda}$
and so
$\left|\frac{d^{+}(v)}{2}-\lambda\right|\leq\left|\frac{d^{+}(v)}{2}-t+1\right|.$
Therefore
$0\leq\left(\frac{d^{+}(v)}{2}-t+1\right)^{2}-\left(\frac{d^{+}(v)}{2}-\lambda\right)^{2}=(d^{+}(v)-t-\lambda+1)(\lambda-t+1).$
The “furthermore” part is obvious. ∎
Lemma 3.2.
Let $D$ be a $(t,\lambda)$-liking digraph. If $t\geq\lambda+1$ or $d^{+}(v)=t+\lambda-1$ for each vertex $v$, then $d^{+}(v)=d^{-}(v)$ for every vertex $v$ in $D$.
Proof.
Assume $t\geq\lambda+1$. To the contrary, suppose that there is a vertex whose outdegree and indegree are different. Then there is a vertex $v$ such that $d^{+}(v)<d^{-}(v)$. Thus
$0<(d^{+}(v)-t-\lambda+1)(\lambda-t+1)$
by the “furthermore” part of Lemma 3.1. Then $d^{+}(v)\neq t+\lambda-1$. Since $d^{+}(v)\geq t+\lambda-1$ by Lemma 2.1, $\lambda+1>t$. ∎
Lemma 3.3.
Let $D$ be a $(t,\lambda)$-liking digraph. If $d^{+}(v)=t+\lambda-1$ for each vertex $v$, then $D$ is the complete digraph on $t+\lambda$ vertices.
Proof.
Suppose that $d^{+}(v)=t+\lambda-1$ for each vertex $v$. Then $d^{-}(v)=t+\lambda-1$ for each vertex $v$ by Lemma 3.2. Therefore
$\sum_{v\in V(D)}{d^{-}(v)\choose t}=\sum_{v\in V(D)}{t+\lambda-1\choose t}=|V(D)|{t+\lambda-1\choose t}.$
Now, by Lemma 2.2(1),
$\frac{t}{|V(D)|}{|V(D)|\choose t}=\frac{t}{\lambda}{t+\lambda-1\choose t}.$
We note that
(${|V(D)|-1\choose t-1}=\frac{t}{|V(D)|}{|V(D)|\choose t}\quad\text{and}\quad\frac{t}{\lambda}{t+\lambda-1\choose t}={t+\lambda-1\choose t-1}.$
Thus
(${|V(D)|-1\choose t-1}={t+\lambda-1\choose t-1}.$
and so $|V(D)|=t+\lambda$. Since $d^{+}(v)=t+\lambda-1$ for each vertex $v$, $N^{+}(v)=V(D)\setminus\{v\}$. Therefore $D\cong\overleftrightarrow{K}_{t+\lambda}$. ∎
Now, we prove Theorem 1.1.
Proof of Theorem 1.1.
Suppose that $D$ is a $(t,\lambda)$-liking digraph with $t\geq\lambda+2$ and take a vertex $v$. By Lemma 3.2,
$d^{+}(v)=d^{-}(v).$
Then, by Lemma 3.1,
$0\leq(d^{+}(v)-t-\lambda+1)(\lambda-t+1).$
Since $t\geq\lambda+2$ by the hypothesis, $\lambda-t+1<0$ and so $d^{+}(v)\leq t+\lambda-1$. Thus, by Lemma 2.1, $d^{+}(v)=t+\lambda-1$. Since $v$ was arbitrarily chosen, $d^{+}(u)=t+\lambda-1$ for each vertex $u$. Therefore $D$ is isomorphic to $\overleftrightarrow{K}_{t+\lambda}$ by Lemma 3.3. ∎
To prove Theorem 1.2, we go further to derive more results. For a vertex set $S$ in a digraph $D$, we denote the set of common out-neighbors (resp. in-neighbors) of $S$ in $D$ by $CN_{D}^{+}(S)$ (resp. $CN_{D}^{-}(S)$), that is,
$CN_{D}^{+}(S):=\bigcap_{u\in S}N_{D}^{+}(u)\quad\text{and}\quad CN_{D}^{-}(S):=\bigcap_{u\in S}N_{D}^{-}(u).$
If no confusion is likely, we write $CN^{+}(S)$ and $CN^{-}(S)$ instead of $CN_{D}^{+}(S)$ and $CN_{D}^{-}(S)$, respectively.
Lemma 3.4.
Let $D$ be a $(t,\lambda)$-liking digraph. If $D$ is a $(t-1,\lambda+1)$-liking digraph, then $D$ is the complete digraph on $t+\lambda$ vertices.
Proof.
Suppose that $D$ is a $(t-1,\lambda+1)$-liking digraph. We first claim the following.
Claim A.
Let $S$ be a set of $t-1$ vertices in $D$. Then the subdigraph of $D$ induced by the common out-neighbors of the vertices in $S$ is complete.
Proof of Claim A.
Since $D$ is a $(t-1,\lambda+1)$-liking digraph and $|S|=t-1$,
$\left|CN^{+}(S)\right|=\lambda+1.$ (1)
We take a common out-neighbor $v$ of the vertices in $S$. Then $v\not\in S$ and so $|S\cup\{v\}|=t$. Thus, since $D$ is a $(t,\lambda)$-liking digraph,
$\left|CN^{+}(S)\cap N^{+}(v)\right|=\lambda.$
By (1),
$CN^{+}(S)\cap N^{+}(v)=CN^{+}(S)-\{v\}.$
Since $CN^{+}(S)\cap N^{+}(v)\subseteq N^{+}(v)$, we have
$CN^{+}(S)-\{v\}\subseteq N^{+}(v).$
Therefore the vertices in $CN^{+}(S)$ other than $v$ are out-neighbors of $v$. Since $v$ was arbitrarily chosen from $CN^{+}(S)$, the subdigraph of $D$ induced by $CN^{+}(S)$ is complete. ∎
Let $U$ be a maximum subset of $V(D)$ inducing a complete subdigraph in $D$. It suffices to show that $U=V(D)$. By Claim A, $|U|\geq\lambda+1$. We may assume $t\leq\lambda+1$, or else we are done by Theorem 1.1. Thus $|U|\geq t$.
To the contrary, suppose that there is a subset $S$ of $U$ such that $|S|=t-1$ and $CN^{+}(S)\not\subseteq U$. Take
$x\in CN^{+}(S)-U.$
By the choice of $U$, $U-S\subseteq CN^{+}(S)$ and so $x\in CN^{+}(U-S)$ by Claim A. Since $x\in CN^{+}(S)$, $x\in CN^{+}(U)$. Now, we take $y\in U$. Since $|U|\geq t$, there is a subset $S^{\prime}$ of $U-\{y\}$ such that $|S^{\prime}|=t-1$. Then, by the choice of $U$, $U-S^{\prime}\subseteq CN^{+}(S^{\prime})$. Since $x\in CN^{+}(U)$, $x\in CN^{+}(S^{\prime})$ and so, by Claim A, $x\in CN^{-}(U-S^{\prime})$. Thus $x\in CN^{-}(U-S^{\prime})\subseteq N^{-}(y)$. Since $y$ was arbitrarily chosen in $U$, $x\in CN^{-}(U)$. Then $U\cup\{x\}$ induces a complete subdigraph in $D$, which contradicts the choice of $U$. Therefore we may conclude that for every subset $S$ of $U$ with $|S|=t-1$,
$CN^{+}(S)\subseteq U.$
Hence $|U|=t+\lambda$ since $D$ is a $(t-1,\lambda+1)$-liking digraph.
Suppose for contradiction that there is a vertex $x$ in $V(D)-U$. Take a subset $S\in U$ with $|S|=t-1$. Then $CN^{+}(S)=U-S$ and so $CN^{+}(S\cup\{x\})\subseteq U-S$. Since $|S\cup\{x\}|=t$, $|CN^{+}(S\cup\{x\})|=\lambda$. Since $|U-S|=\lambda+1$, there is a unique vertex in $U-S$ that is not an out-neighbor of $x$. If there were at least two vertices $y$ and $z$ in $U$ such that $y\not\in N^{+}(x)$ and $z\not\in N^{+}(x)$, then selecting $S\subseteq U-\{y,z\}$ would be a contradiction. If there were only one vertex $y$ in $U$ such that $y\not\in N^{+}(x)$, then selecting $S\in U$ with $y\in S$ would be a contradiction. Therefore, in any cases, we have reached a contradiction. Hence $U=V(D)$ and so $D$ is isomorphic to $\overleftrightarrow{K}_{t+\lambda}$. ∎
Lemma 3.5.
Let $D$ be a $(t,\lambda)$-liking digraph for some integers $(t,\lambda)\neq(2,1)$. If there exists a vertex $v$ such that $N^{+}(v)=V(D)-\{v\}$, then $D$ is the complete digraph on $t+\lambda$ vertices.
Proof.
Suppose that there exists a vertex $v$ such that
$N^{+}_{D}(v)=V(D)-\{v\}.$
Then $D-v$ is a $(t-1,\lambda)$-liking digraph by Proposition 2.3. As long as $v$ is an out-neighbor of every other vertex, i.e.
$N^{-}_{D}(v)=V(D)-\{v\},$
$D-v$ is a $(t,\lambda-1)$-liking digraph. Then $D-v$ is isomorphic to $\overleftrightarrow{K}_{t+\lambda-1}$ by Lemma 3.4. Therefore $D$ is isomorphic to $\overleftrightarrow{K}_{t+\lambda}$. Thus it is sufficient to show $N^{-}_{D}(v)=V(D)-\{v\}$.
To this end, we suppose, to the contrary, that there is a vertex $w$ such that $w\neq v$ and $w\not\to v$. Since $|V(D)|\geq t+\lambda\geq t+1$ by Lemma 2.1, there is a subset $T$ of $V(D)-\{v\}$ such that $|T|=t-1$ and $w\in T$. Then $|CN_{D}^{+}(T)|\geq\lambda+1$ by Proposition 2.4. Since $w\not\to v$, $v\notin CN_{D}^{+}(T)$. Therefore
$CN_{D}^{+}(T)\subseteq V(D)-\{v\}=N_{D}^{+}(v)$
and so
$\left|CN_{D}^{+}(T\cup\{v\})\right|=\left|N_{D}^{+}(v)\cap CN_{D}^{+}(T)\right|=\left|CN_{D}^{+}(T)\right|\geq\lambda+1.$
Thus the vertices in $\{v\}\cup T$ have at least $\lambda+1$ common out-neighbors. However, |$|\{v\}\cup T|=t$, which contradicts the hypothesis that $D$ is a $(t,\lambda)$-liking digraph. Thus $v$ is an out-neighbor of every other vertex in $D$. ∎
Now, we are ready to prove Theorem 1.2.
Proof of Theorem 1.2.
It is obvious that $(a)$ implies each of $(b)$, $(c)$, $(d)$, and $(e)$. By Lemma 3.4, $(b)$ implies $(a)$. By Lemma 3.3, $(c)$ implies $(a)$. Thus $(a)$, $(b)$, and $(c)$ are equivalent. If $(t,\lambda)\neq(2,1)$, then Lemma 3.5 guarantees that $(d)$ implies $(a)$. If $t\geq 3$, then Proposition 2.7 guarantees that $(e)$ implies $(a)$. Thus the “furthermore” part is true. This completes the proof. ∎
4 Concluding Remarks
According to [4] and [16], a $(t,\lambda)$-friendship graph is the complete graph on $t+\lambda$ vertices if $t\geq 3$. This result is partially established for $(t,\lambda)$-liking digraphs with $t\geq\lambda+2$ by our result, Theorem 1.1. We would like to know if this digraph version holds in general.
Conjecture 4.1.
The complete digraph on $t+\lambda$ vertices is the only $(t,\lambda)$-liking digraph if $t\geq 3$.
Meanwhile, according to [3], a $(2,\lambda)$-friendship graph is regular if $\lambda\geq 2$. Together with the above observation, it is true that a $(t,\lambda)$-friendship graph is regular for any integers $t\geq 2$ and $\lambda\geq 1$ with $(t,\lambda)\neq(2,1)$. It is natural to question whether this result also holds for a $(t,\lambda)$-liking digraph for any integers $t\geq 2$ and $\lambda\geq 1$ with $(t,\lambda)\neq(2,1)$. For the case $t\geq 3$, the answer to the given question immediately follows the equivalence of (a) and (e) in Theorem 1.2.
Corollary 4.2.
If $t\geq 3$, then the complete digraph on $t+\lambda$ vertices is the only diregular $(t,\lambda)$-liking digraph.
Yet, this distinction breaks down at $t=2$, as illustrated by our explicit $(2,2)$-liking digraph example that is not diregular (see Figure 1).
Figure 1: A $(2,2)$-liking digraph which is not diregular
5 Acknowledgement
This work was supported by Science Research Center Program through the National Research Foundation of Korea(NRF) grant funded by the Korean Government (MSIT)(NRF-2022R1A2C1009648). Hojin Chu was partially supported by a KIAS individual Grant (CG101801) at Korea Institute for advanced Study.
The authors are grateful to Taehee Hong and Homoon Ryu for helping us code the exhaustive search and find a counterexample.